Find the equation of the line which passes through the point (1,-2),
and has slope = 3

Answer by Edwin McCravy(20063)

   y - y1 = m(x - x1)

Substitute (x1,y1) = (1,-2), m = 3

y - (-2) = 3(x - 1)

   y + 2 = 3x - 3

       y = 3x - 5

Edwin